Write an equation and solve. You eat 4 pieces of a large pizza. THere is 2/3 of a pizza left. How many pieces make a large pizza?
Hello I am having trouble putting this problem into an equation, I know that the final answer is 12, I just don't know how to put it into an equation, please help me, thanks!

Answer by stanbon(75887)

Write an equation and solve. You eat 4 pieces of a large pizza. THere is 2/3 of a pizza left. How many pieces make a large pizza?
Hello I am having trouble putting this problem into an equation, I know that the final answer is 12, I just don't know how to put it into an equation, please help me, thanks!
-----
Let the number of pieces in the whole pizza be "x"::
Equation:
x-4 = (2/3)x
-----
(1/3)x = 4
-----
x = 12 pieces
----------
Cheers,
Stan H.